# Meta Co-Training: Two Views are Better than One

### v2

@@ 4.2 Meta Co-Training
6:     // Predict soft pseudo labels
@@
9:     // Sample the pseudo labels from the discrete distribution over the classes
@@
12:     // MPL student update applied to both models
@@
18:     // MPL teacher update applied to both models
@@
Algorithm 1 Meta Co-Training (Section 3.1 has notation)
At each step $t$ of the algorithm each model is first updated based on the pseudo labels the other has provided on a batch of unlabeled data (Lines 16-17). Then each model is updated to provide labels that encourage the other to predict more correctly on the labeled set based on the performance of the other (Lines 23-24). We refer the reader to Section 3.2 for a brief explanation as to why these updates make sense, or to [59] for a more complete discussion of the MPL approach. Geometrically, we can understand this as updating the parameters in the direction of increasing confidence on our pseudo labels if those labels helped the other model perform better on the labeled set, or in the opposite direction if it hurt the other model’s performance on the labeled set. Note that contrary to traditional co-training, no new pseudo-labeled data are added to the labeled set, neither are any data removed from the unlabeled set after each step.
5 Experiments
@@
Towards evaluating the proposed methods of view construction and co-training we first constructed views of the data utilizing five different learned representations. Our primary comparisons were on the widely-used ImageNet [31] classification task; we provide experimental results on additional image classification tasks. To produce the views used to train the classifiers during co-training and meta co-training we used the embedding space of five representation learning architectures: The Masked Autoencoder (MAE) [42], DINOv2 [58], SwAV [19], EsViT [49], and CLIP [61]. We selected the models which produce the views as they have been learned in an unsupervised way, have been made available by the authors of their respective papers for use in PyTorch, and have been shown to produce representations that are appropriate for ImageNet classification.
@@
On the Sufficiency of the Views.   Motivated by the analysis and experiments of [34] we made an effort to verify the sufficiency and independence of the representations learned by the aforementioned models. Sufficiency is fairly easy to verify. If we choose a reasonable sufficiency threshold for the task, for ImageNet say $75\%$ top-1 accuracy, then we can train simple models on the views of the dataset we have constructed and provide examples of functions which demonstrate the satisfaction of the property. We tested a single linear layer with softmax output (Table 1), and a simple 3-layer multi-layer perceptron (MLP) with 1024 neurons per layer and dense skip connections (previous layer outputs were concatenated to input for subsequent layers) (Table 2) to provide a lower bound on the sufficiency of the views on the standard subsets of the ImageNet labels for semi-supervised classification.
@@
Model 1% 10%
MAE 23.4 48.5
DINOv2 78.4 82.7
SwAV 12.5 32.6
EsViT 71.3 75.8
CLIP 75.2 80.9
From these experiments we can see that if we were to pick the $75\%$ top-1 accuracy threshold, which we believe indicates considerable skill without presenting too high of a bar, then CLIP and DINOv2 provide sufficient views for both the $1\%$ and $10\%$ subsets.
On the Independence of the Views.   To test the independence of the views generated by the representation learning models, we trained an identical MLP architecture to predict each view from each other view; Table 3 presents our findings, explained below.
MAE DINOv2 SwAV EsViT CLIP
MAE - 0.139 0.142 0.137 0.129
DINOv2 0.110 - 0.132 0.135 0.130
SwAV 0.116 0.147 - 0.137 0.126
EsViT 0.112 0.140 0.116 - 0.135
CLIP 0.110 0.146 0.136 0.156 -
@@
Co-Training Experiments on ImageNet.   Having constructed at least two strong views of the data, and suspecting that these views are independent we hypothesize that co-training will work on this data. In Table 3(a) we show the results of applying the standard co-training algorithm [12]. At each iteration of co-training the model made a prediction for all of the instances in $U$, following (4). We assigned labels to the 10% of the total original unlabeled data with most confident predictions for both models. When confident predictions conflicted on examples, we returned those instances to the unlabeled set, otherwise they entered the labeled set of the other model with the assigned pseudo-label.
@@
Comparing co-training accuracy on ImageNet 10% (resp. 1%) shown in Table 3(a) compared to MLP accuracy shown in Table 2, we observe that co-training top-1 accuracy is on average 18.7% (resp. 16.8%) higher than MLP accuracy. Moreover, co-training top-1 accuracy is 2.4% (resp. 1.7%) higher overall. In other words, co-training resulted in a significant increase in prediction accuracy but not after any iterations of co-training have occurred; see Figures 1 and 2. Furthermore, predicting according to (4) yielded better accuracy than the predictions of each individual model. In addition, in our experiments on ImageNet, as co-training iterations proceeded top-1 accuracy decreased. Later we show that this was not always the case (see Figure 9 in the appendix). The decrease was less pronounced when the views perform at a similar level (see Figure 3). Despite the poor performance of the pseudo-labeling during co-training, just computing the re-normalized product of the two best performing views performed as well as the previous SotA method for semi-supervised classification on the ImageNet-10% task (see Co-Training in Table 5). The performance of co-training as a method was rather disappointing in that the algorithm is unable to leverage pseudo-labels to improve the accuracy on this task; however, the results do show that our view-construction method has merit. Given the difficulty of translating between the views, and the fact that the accuracy yielded by (4) is greater than its constituent views, we believe the poor performance of the pseudo-labeling step in co-training was due to the insufficiency of at least one view rather than view-dependence.
Figure 1: Top-1 accuracy of co-training iterations for the ImageNet 10% dataset. The average over all 10 pairs, the best pair, and the worst pair at each iteration are shown.
Figure 2: Top-1 accuracy of co-training iterations for the ImageNet 1% dataset. The average over all 10 pairs, the best pair, and the worst pair at each iteration are shown.
@@
Figure 3: Top-1 accuracy of co-training iterations on the CLIP and DINOv2 views for the ImageNet 10% dataset.
@@
Figure 4: Top-1 accuracy of co-training iterations on the CLIP and DINOv2 views for the ImageNet 1% dataset.
DINOv2 SwAV EsViT CLIP
MAE 81.8 (75.5) 30.5 (11.4) 77.1 (66.2) 78.8 (66.8)
DINOv2 78.5 (74.5) 83.3 (78.9) 85.1 (80.1)
SwAV 70.6 ( 64.9) 75.5 (67.4)
EsViT 82.3 ( 76.9)
(a) Co-Training
DINOv2 SwAV EsViT CLIP
MAE 81.2 (73.8) 37.6 (8.1) 73.6 ( 63.8) 78.6 (63.5)
DINOv2 77.3 ( 70.7) 83.4 (79.2) 85.2 (80.7)
SwAV 74.4 (67.3) 77.1 (67.4)
EsViT 82.4 (77.5)
(b) Meta Co-Training
Table 4: Co-training and meta co-training top-1 accuracy of view combinations on 10% (1%) of ImageNet labels. As co-training does not depend on the order of the views, we show only upper-diagonal entries of the pairwise comparison. Diagonal entries would correspond to views which are trivially dependent. Note that all $\binom{5}{2}=10$ different combinations of views are shown in each table.
@@
MCT (MLP)* - ours MCT 80.7 85.8
@@
Meta Co-Training Experiments on ImageNet.   To draw fair comparisons we fix the model architecture and view set from our experiments on co-training. As in [59] a supervised loss is optimized jointly with the MPL loss. We found that beginning with a warmup period in which the model trained in a strictly supervised way expedited training. Unlike MPL, we did not use Unsupervised Data Augmentation [75], because we embedded the dataset only once, and due to the nature of the pretext tasks that were used to generate the views we did not believe data augmentation would have much of an effect. Details of our training recipe appear in Table 10.
@@
Comparing meta co-training accuracy on ImageNet 10% (resp. 1%) shown in Table 3(b) compared to MLP accuracy shown in Table 3(a), we observe that meta co-training top-1 accuracy yielded a difference on average 0.7% (resp. -0.6%) compared to co-training top-1 accuracy. Moreover, meta co-training top-1 accuracy is 0.1% (resp. 0.6%) higher overall. Note that on all view pairs in which co-training beats meta co-training, with the exception of the $10\%$ (EsViT, MAE) pair, one of the views alone was better than co-training. The single DINOv2 view outperforms all such view pairs. These are all cases in which one of the views was significantly weaker than the other. For the experiments in which we had two strong (e.g., not SwAV or MAE) views, meta co-training always had better performance than co-training. Finally, we take the four views which had the greatest performance and constructed two views out of them by concatenating them together. We constructed one view CLIP |||| EsViT and the second DINOv2 |||| SwAV and measured the performance of co-training and meta co-training on these views in Table 6. Interestingly, co-training does not benefit from these larger views; see Table 6.
@@
MCT 80.5 85.8
Table 6: Meta co-training and co-training evaluated on the ImageNet dataset using the CLIP |||| EsViT and DINOv2 |||| SwAV views.
The additional information may have allowed the individual models to overfit their training data quickly. During meta co-training we observed that after one model reaches 100% training accuracy, its validation accuracy can still improve by learning to label for the other model. Possibly due to the rapid overfitting due to the increased view size (and consequently increased parameter count) the 1% split did not show improvement for either model.

### v5

@@ 3 Proposed Method: Meta Co-Training
6:     // Predict soft pseudo-labels
@@
9:     // Sample the pseudo-labels from the discrete distribution over the classes
@@
12:     // MCT student update applied to both models
@@
18:     // MCT teacher update applied to both models
@@
Algorithm Description.
Algorithm 1 presents pseudocode for the algorithm. The function SampleBatch, is sampling a batch from the dataset. The function getOtherView returns the complementary view of the first argument; the second argument clarifies which view should be returned (“2” in line 5). At each step $t$ of the algorithm each model is first updated based on the pseudo-labels the other has provided on a batch of unlabeled data (lines 13-14). Then each model is updated to provide labels that encourage the other to predict more correctly on the labeled set based on the performance of the other (lines 24-25). Further discussion of these update rules can be found in Appendix A.
Each student model evaluates the performance of their new weights on separate labeled batches. This performance provides feedback to the teacher model. Geometrically, we can understand this as updating the parameters in the direction of increasing the teacher’s confidence on pseudo-labels, assuming those pseudo-labels helped the student model perform better on the labeled set. Along these lines, the teacher parameters are updated in the opposite direction if it hurt the student’s performance on the labeled set.
In general MCT does not require any preexisting method of view construction to be applied. If a dataset naturally provides two views, then there is no need to construct additional views. One can use their favorite pre-trained representation or none at all and then apply MCT as described above. Additionally we can relax the assumption that there is correspondence between the labeled examples for each view. If it happens to be the case that it is easy to collect unlabeled examples from each view which correspond to the same label but existing labeled sets for each view do not, then MCT can still be applied. This is because the labeled data used to evaluate the change in the performance of the student does not need to have corresponding instances in the teacher’s view. In all of our experiments in this work we use preexisting foundation models to construct views for our co-training methods, and thus in all cases there exist two views for each labeled example.
In Algorithm 1 it is illustrative to consider all versions of the parameters within a step and all of their gradients to exist simultaneously in memory. This is undesirable in practice given the large size of common neural networks and their gradients. Instead, we compute $\nabla_{\theta_{2}}\ell(PL^{(2)},\hat{y}^{(2)})$ and $\nabla_{\theta_{1}}\ell(PL^{(1)},\hat{y}^{(1)})$ (used in lines 24-25) first, then we compute (and apply) $\nabla_{\theta_{1}}\ell(PL^{(2)},\hat{y}^{(1)})$ and $\nabla_{\theta_{2}}\ell(PL^{(1)},\hat{y}^{(2)})$ (used first in lines 13-14 and reused in lines 21-22), then we compute $\nabla_{\theta^{\prime}_{1}}\ell(Y_{t}^{(1)},\hat{y}^{\prime(1)})^{T}$ and $\nabla_{\theta^{\prime}_{2}}\ell(Y_{t}^{(2)},\hat{y}^{\prime(2)})^{T}$ which we subsequently use to compute $h^{(1)}$ and $h^{(2)}$. We finish the iteration with lines 24-25 using the saved gradient from the first step. With this approach we only have to compute the forward pass and gradient for the student update once, and we only keep one version of the model weights in memory at a time.
Relation to Ensemble Methods.
As our approach utilizes multiple models which collaborate during prediction, we acknowledge that it bears similarity to the family of ensemble methods. Typically, co-trained models are evaluated against other SSL approaches. In our evaluation we compare our semi-supervised method to supervised deep ensembles to illustrate the impact of MCT versus ensemble methods. We show that when MCT works well, the performance benefit cannot be attributed to simply utilizing multiple models.
4 Experimental Evaluation
@@
Towards evaluating our proposed method we compare to other self-supervised and semi-supervised learning algorithms using the ImageNet [17] classification benchmark, as well as Flowers102 [41], Food101 [6], FGVCAircraft [39], iNaturalist [29], and iNaturalist 2021 [28] datasets. To produce the views used to train the classifiers during CT and MCT we used the embedding space of five representation learning architectures: The Masked Autoencoder (MAE) [26], DINOv2 [42], SwAV [10], EsViT [34], and CLIP [45]. We selected the models which produce the views as they have been learned in an unsupervised way, have been made available by the authors of their respective papers for use in PyTorch, and have been shown to produce representations that are appropriate for computer vision classification tasks. Hyperparameters used during training are included in Table 11 in the appendix for CT, MCT, and deep ensembles.
Table 1 provides a description of the different datasets used for training and testing, as well as the number of classes for each dataset. Note that in the 10% experiments on the Flowers102 dataset each class has only 1 label. All datasets are approximately class-balanced with the exception of the iNaturalist dataset. In all cases we maintain the original class distribution when sampling subsets. For ImageNet we use the split published with the SimCLRv2 [13] repository. All subsets are created with seeded randomness of a common seed (13) which ensures a fair comparison between methods.
Table 1: Characteristics of datasets used.
Dataset #train #test #classes
FGVCAircraft $3$ $3$ 100
Flowers102 1020 6149 102
Food101 75,750 25,250 101
iNaturalist 175,489 29,083 1010
ImageNet 1,281,167 50,000 1000
4.1 Experimental Evaluation on ImageNet
As we discussed in Section 2, CT relies on two assumptions about the sufficiency and independence of the two views in our data. We conducted experiments to verify these assumptions. Below we give a summary of our results.
On the Sufficiency of the Views.
@@ 4.1 Experimental Evaluation on ImageNet
Sufficiency is fairly easy to verify. If we choose a reasonable sufficiency threshold for the task, for ImageNet say close or above $75\%$ top-1 accuracy, then we can train simple models on the views of the dataset we have constructed and provide examples of functions which demonstrate the satisfaction of the property. We tested a single linear layer with softmax output to provide a lower bound on the sufficiency of the views on the standard subsets of the ImageNet labels for semi-supervised classification.
A simple 3-layer multi-layer perceptron (MLP) with 1024 neurons per layer was evaluated (Table 2). These were the same models as those used for CT and MCT. From these experiments, whose results are shown in Table 2 we can see that if we were to pick a threshold for top-1 accuracy around $75\%$, then CLIP and DINOv2 provide sufficient views for both the $1\%$ and $10\%$ subsets.
Model 1% 10%
MAE 23.4 48.5
DINOv2 78.4 82.7
SwAV 12.5 32.6
EsViT 71.3 75.8
CLIP 75.2 80.9
On the Independence of the Views.
To test the independence of the views generated by the representation learning models, we trained an identical MLP architecture to predict each view from each other view; Table 3 presents our findings, explained below.
MAE DINOv2 SwAV EsViT CLIP
MAE - 0.139 0.142 0.137 0.129
DINOv2 0.110 - 0.132 0.135 0.130
SwAV 0.116 0.147 - 0.137 0.126
EsViT 0.112 0.140 0.116 - 0.135
CLIP 0.110 0.146 0.136 0.156 -
@@
Having constructed at least two strong views of the data, and suspecting that these views are independent we hypothesize that MCT and CT will work on this data.
Co-Training Baseline Experiments.
@@
At each iteration of CT the models made predictions for all instances in $U$. We assigned labels to the 10% of the most confident predictions in $U$ for both models. When confident predictions conflicted on examples, we returned those instances to the unlabeled set, otherwise their complementary views were entered into the labeled set of the other model with the assigned pseudo-label.
For CT, joint predictions yielded better accuracy than the predictions of each individual model. In our experiments on ImageNet, as CT iterations proceeded, top-1 accuracy decreased. Later we show that this was not always the case (see Figure 4). The decrease was less pronounced when the views perform at a similar level; see Figure 2. In contrast MCT does not exhibit performance degradation after warmup. Figures 2 and 3 show results on ImageNet 10%. We see that while the CT performance decreases after the supervised warmup, MCT performance increases.
@@
Despite the poor performance of the pseudo-labeling during CT, using the two best performing views CT performed as well as the previous SoTA method for SSL classification on the ImageNet-10%; see Co-Training in Table 5. CT as a method was rather disappointing in that the algorithm is unable to leverage pseudo-labels to improve the accuracy on this task. Given the difficulty of translating between the views, and the fact that the accuracy yielded by the joint prediction is greater than its constituent views, we believe the poor performance of the pseudo-labeling step in CT was due to imbalanced information content between views rather than view-dependence.
@@
Figure 2: Top-1 accuracy of CT iterations on the CLIP and DINOv2 views for the ImageNet 10% dataset.
@@
Figure 3: MCT using the CLIP and DINOv2 views as a function of the training step. Models are trained on 10% of the ImageNet labels.
@@
[51] UDA (ResNet50) UDA - 68.78
[49] FixMatch (ResNet-50) FixMatch - 71.46
[43] MPL (EfficientNet-B6-Wide) MPL - 73.9
@@
[32] Deep Ensemble * Deep Ensemble 80.0 84.3
Meta Co-Training (MLP)* - ours Meta Co-Training 80.7 85.8
Meta Co-Training Experiments.
@@
To draw fair comparisons, we fix the model architecture and view set from our experiments on CT. As in MPL [43] a supervised loss is optimized jointly with a loss on pseudo-labels. We found that beginning with a warmup period in which the models for each view are trained in a strictly supervised way expedited training. This warmup period occurs for the same number of updates as the first training phase of CT, so each method starts with the same number of supervised updates before pseudo-labeling. Unlike MPL, we did not use Unsupervised Data Augmentation [51], because we embedded the dataset only once, and using self-supervised models to generate the views we did not believe data augmentation would have much of an effect. Details of our training recipe are given in Table 11.
Table 6 shows the results of MCT on all of the possible combinations of views we used. With one exception, the better the performance of the views, the better MCT performs compared to CT. In nearly all cases in which both views are strong, MCT performs better than CT. These are also the cases with the highest performance overall. Comparing CT accuracy on ImageNet 10% (resp. 1%) shown in Table 6(a) compared to MLP accuracy shown in Table 2, we observe that CT top-1 accuracy is on average 18.7% (resp. 16.8%) higher than MLP accuracy. Moreover, the best CT top-1 accuracy is 2.4% (resp. 1.7%) higher compared to the best top-1 MLP accuracy. In Table 6(c) we show the performance of an ensemble of models trained on concatenated view pairs but we provide more information on the comparison between MCT and deep ensembles in a separate paragraph below.
Table 6: CT, MCT, and ensemble top-1 accuracy of view combinations on 10% (1%) of ImageNet labels. As CT and MCT do not depend on the order of the views, we show only upper-diagonal entries of the pairwise comparison. All $\binom{5}{2}=10$ different combinations of views are shown in each table. In the case of the MLP ensemble the order of the pairs of views also does not matter. These views are concatenated in order to form a larger unified view and hence different order of the views would correspond to different order of the attributes in the embeddings, which does not change their information content.
(a) Co-Training
DINOv2 SwAV EsViT CLIP
MAE 81.8 (75.5) 30.5 (11.4) 77.1 (66.2) 78.8 (66.8)
DINOv2 78.5 (74.5) 83.3 (78.9) 85.1 (80.1)
SwAV 70.6 ( 64.9) 75.5 (67.4)
EsViT 82.3 ( 76.9)
(b) Meta Co-Training
DINOv2 SwAV EsViT CLIP
MAE 81.2 (73.8) 37.6 (8.1) 73.6 ( 63.8) 78.6 (63.5)
DINOv2 77.3 ( 70.7) 83.4 (79.2) 85.2 (80.7)
SwAV 74.4 (67.3) 77.1 (67.4)
EsViT 82.4 (77.5)
(c) MLP Ensemble
DINOv2 SwAV EsViT CLIP
MAE 83.0 (78.9) 38.1 (17.9) 75.0 ( 72.1) 79.4 (76.1)
DINOv2 83.7 ( 79.1) 81.2 (78.5) 84.2 (80.0)
SwAV 74.4 (72.5) 81.2 (76.8)
EsViT 78.7 (73.5)
MCT and CT are compared on ImageNet 10% (resp., 1%) in Table 5. We observe that MCT top-1 accuracy is 0.7% (resp., 0.6%) higher overall. An ablation study for the views is given in Table 6. On all view pairs in which MCT or CT provide no benefit over a single view are cases in which one of the views was significantly weaker than the other. This makes sense as one model has nothing to learn from the other. For the experiments in which we had two strong views, MCT always outperforms CT.
Comparison to Ensemble Methods.
Ensemble methods are common in practice to boost the performance of supervised models. To show that our method provides benefit outside of just adding an additional model, we compare to a Deep Ensemble [32] of five models for each individual view. In Table 7 we show the performance of ensembles trained on the individual views. We show the performance of an ensemble trained on each view pair concatenated in Table 6(c). We show the performance of an ensemble trained on the concatenation of all four views we used to achieve our best performing MCT experiment in Table 8.
In some cases the CT models perform better than the ensemble of models which all have the same input for each sample (the concatenation of the CT views). We believe that CT before pseudo-labeling (which is essentially an ensemble of two different views) out-performs the deep ensembles because the CT models are fewer but they are more diverse. The CT models are more diverse as they are trained on different views, while the ensemble models are trained on the concatenation of the views. The concatenation has the same information but may not produce meaningfully different models in an ensemble. Ensuring diversity of ensemble members is an important problem for ensembling methods, so it seems likely that a less diverse ensemble would perform worse. If this is indeed the case, then it would support our hypothesis that CT is failing to improve after pseudo-labeling due to an imbalance in the sufficiency of views and not a lack of independence. MCT provides performance gains over CT and the deep ensembles, so we can be confident that these gains cannot be attributed solely to the use of multiple models.
Table 7: MLP ensemble performance on individual views. Top-1 accuracy on different subsets of the ImageNet data are shown.
Model 1% 10%
MAE 1.7 3.8
DINOv2 79.2 82.8
SwAV 19.7 40.3
EsViT 72.1 75.0
CLIP 76.4 80.8
4.2 Stronger Views by Concatenation
@@
Finally, we take the four views which had the greatest performance and constructed two views out of them by concatenating them together. We constructed one view CLIP |||| EsViT and the second DINOv2 |||| SwAV and measured the performance of CT and MCT on these views in Table 8. Interestingly, CT does not benefit from these larger views; see Table 8.
Table 8: Co-training and MCT evaluated on the ImageNet dataset using the CLIP |||| EsViT and DINOv2 |||| SwAV views. The Deep Ensemble is evaluated on the concatenation of all four views.
@@
Meta Co-Training 80.5 85.8
The additional information may have allowed the individual models to overfit their training data quickly. During MCT we observed that after one model reaches 100% training accuracy, its validation accuracy can still improve by learning to label for the other model. Possibly due to the rapid overfitting due to the increased view size (and consequently increased parameter count) the 1% split did not show improvement for either model.
4.3 Experiments on Additional Datasets
We compare MCT and CT to existing SoTA approaches that leverage unlabeled data directly during training (e.g., not just part of a pretext task). These experiments support our hypothesis that the primary cause of the failure of CT is that one or more of the views is sufficient to learn a model that achieves high accuracy. The closer to equivalent in performance individual views are, the less performance suffers. When the two views perform similarly and well, the performance improves. Tables 9 and 10 provide more information on these additional datasets. Furthermore, Figure 4 provides an example where CT performs as expected and accuracy increases across CT iterations. In all but one case MCT outperforms the SoTA top-1 accuracy. The main takeaway of these additional experiments, apart from establishing new SoTA results in datasets beyond ImageNet, is that MCT is in general more robust to view imbalance and provides better results than CT. As the performance degradation of co-training beyond the initial warmup period was not something expected, in Appendix D we provide charts similar to the one in Figure 4, but for the other cases of datasets or presence of labels in the training set, thus, giving full details on the results that we obtained for CT.
Table 9: Additional comparisons to REACT. The authors include experiments for zero-shot performance and 10% of available labels. Results shown are for 10% of labels. For Flowers102 this is 1-shot performance.